A man who can throw a stone with a velocity of 20 m/s wishes to hit a target placed on his own level at a distance of 30m. At what angle should be threw the stone?


1
Expert's answer
2021-06-22T05:08:30-0400
"v_x(t)=v_0\cos \alpha"

"x(t)=x_0+v_0\cos \alpha \cdot t"

"v_y(t)=v_0\sin \alpha-gt"

"y(t)=y_0+v_0\sin \alpha\cdot t-\dfrac{gt^2}{2}"



Let "x_0=x(0)=0, y_0=y(0)=0"

Find the time when "y=0"


"0+v_0\sin \alpha\cdot t-\dfrac{gt^2}{2}=0"

"t_1=0, t_2=\dfrac{2v_0\cdot \sin \alpha}{g}"

Then


"x(t_2)=0+v_0\cos \alpha \cdot (\dfrac{2v_0\cdot \sin \alpha}{g})=30\ m"

"\sin 2\alpha=\dfrac{30\ m\cdot g}{2v_0^2}"

"v_0=20\ m/s, g=9.81\ m/s^2"



"\sin 2\alpha=\dfrac{30\ m\cdot 9.81\ m/s^2}{2(20\ m/s)^2}=0.367875"

Since "0\degree\leq\alpha \leq90\degree"

"2\alpha=\sin^{-1}(0.367875)"

or


"2\alpha=180\degree-\sin^{-1}(0.367875)"

"2\alpha\approx21.5846\degree \text{ or }2\alpha\approx158.4154\degree"

"\alpha\approx10.8\degree \text{ or }\alpha\approx79.2\degree"